# thanks: Corresponding author

### v1

@@ III.4 Radiation damage
Fig. 14 shows the Warren-Cowley short-range order (SRO) parameters [94] (first shell) calculated for all atoms pairs in the equiatomic composition. It is immediately evident that there is preferential ordering that is induced as a consequence of the collision cascades. The negative values mean a tendency for short-range ordering, while positive values indicate the opposite tendency. Al-Al and Ni-Ni pairs show the strongest tendency to separate, while Al-Ni shows strongest tendency to order. This is logical, as it means that the Al-Al and Ni-Ni separate to form Al-Ni ordering. The Cu-Cu, Fe-Fe and Cr-Cr elemental pairs all show a degree of short-range ordering, with Cu also showing significant ordering with Al. Furthermore, we see that the SRO parameters show some degree of saturation although they are still changing well after defect concentration has stabilized (see Fig. 8, around 0.04–0.05 dpa). There are differences in the values of the SRO parameters and ordering of the pairs between the different compositions. However, the overall tendencies are similar. In all cases (that include Al), Al-Al pairs show strongest tendency to separate, while Al-Ni shows strongest tendency for ordering. In the CrCuFeNi composition, Cu-Cu and Fe-Fe pairs show the strongest tendency for ordering. The SRO parameters of the other compositions can be found in the supplementary materials (S4).
III.5 Annealing
@@
To accelerate thermal recombination that cannot be accessed directly within MD time scales, MD annealing runs at elevated temperatures are frequently applied. For tungsten-based alloys, the defect structures in simulation cells of different compositions were shown to exhibit significant differences after annealing, which was explained by the different diffusion rates of interstitial and vacancy-type defects [43]. To investigate whether this effect is also present in the FCC HEAs which are considered in this work, the final irradiated structures were annealed for 1 ns at 800 K. The temperature was chosen to be high enough for defect recombination to take place at an accelerated rate, while still being well below the melting temperatures of the materials. The annealing was performed in the $NPT$ ensemble and consisted of three stages. First, the temperature of the system was raised from 300 K to 800 K during the first 200 ps. Secondly, the system was held at 800 K for the next 600 ps, and finally the system was cooled back to the initial 300 K during the final 200 ps of the simulation. During this process, the vacancy concentration of the systems were tracked. Fig. 15 shows the vacancy concentration as a function of time during annealing. During the simulation, roughly 17 % of the vacancies recombine and anneal out. However, significant differences between the different compositions were not observed, at least at this temperature and time scale. Additionally, the final structures after annealing, while not identical, did not show significant differences between the compositions. The final structures can be found in the supplementary materials (S5).
@@ IV Discussion
Our general-purpose interatomic potential developed for AlCrCuFeNi alloys enables computationally efficient and accurate computational investigations of various alloys. While the resulting tabGAP model does not explicitly include magnetic degrees of freedom, the use of magnetic DFT training data is key to enable simulations of multiphase alloys and phase transitions. This is evident from the Monte Carlo swapping MD simulations, where a dynamic transition from FCC to BCC was observed for the equiatomic composition. Even more strikingly, the MCMD simulations computationally support the VEC stability rule (see Table 1 for VECs), where for alloy compositions with VEC $\geq 8$ a single FCC phase is expected and for $6.87\leq$ VEC $<$ 8 FCC and BCC coexistence is expected [74]. This rule has also found support in experiments [10, 95, 96]. However, the VEC rule might not be well suited for prediction of phases in non-equilibrium rapidly solidified HEAs [15], and BCC phases have been observed with AlCrCuFeNi alloys above the VEC criterion [87]. Furthermore, the addition of large amounts of Al causes significant lattice distortions (due to it having a large atom radius, see Table 1 for $\delta\%$), which makes the close-packed FCC less stable and thus the BCC lattice can accommodate the larger Al atoms more easily in it [97].

### v2

@@ III.5 Radiation damage
Fig. 16 shows the Warren-Cowley short-range order (SRO) parameters [102] (first shell) calculated for all atoms pairs in the equiatomic composition. It is immediately evident that there is preferential ordering that is induced as a consequence of the collision cascades. The negative values mean a tendency for short-range ordering, while positive values indicate the opposite tendency. Al-Al and Ni-Ni pairs show the strongest tendency to separate, while Al-Ni shows strongest tendency to order. This is logical, as it means that the Al-Al and Ni-Ni separate to form Al-Ni ordering. The Cu-Cu, Fe-Fe and Cr-Cr elemental pairs all show a degree of short-range ordering, with Cu also showing significant ordering with Al. Furthermore, we see that the SRO parameters show some degree of saturation although they are still changing well after defect concentration has stabilized (see Fig. 10, around 0.04–0.05 dpa). There are differences in the values of the SRO parameters and ordering of the pairs between the different compositions. However, the overall tendencies are similar. In all cases (that include Al), Al-Al pairs show strongest tendency to separate, while Al-Ni shows strongest tendency for ordering. In the CrCuFeNi composition, Cu-Cu and Fe-Fe pairs show the strongest tendency for ordering. The SRO parameters of the other compositions can be found in the supplementary materials (Fig. S4).
III.6 Annealing
@@
To accelerate thermal recombination that cannot be accessed directly within MD time scales, MD annealing runs at elevated temperatures are frequently applied. For tungsten-based alloys, the defect structures in simulation cells of different compositions were shown to exhibit significant differences after annealing, which was explained by the different diffusion rates of interstitial and vacancy-type defects [47]. To investigate whether this effect is also present in the FCC HEAs which are considered in this work, the final irradiated structures were annealed for 1 ns at 800 K. The temperature was chosen to be high enough for defect recombination to take place at an accelerated rate, while still being well below the melting temperatures of the materials. The annealing was performed in the $NPT$ ensemble and consisted of three stages. First, the temperature of the system was raised from 300 K to 800 K during the first 200 ps. Secondly, the system was held at 800 K for the next 600 ps, and finally the system was cooled back to the initial 300 K during the final 200 ps of the simulation. During this process, the vacancy concentration of the systems were tracked. Fig. 17 shows the vacancy concentration as a function of time during annealing. During the simulation, roughly 17 % of the vacancies recombine and anneal out. However, significant differences between the different compositions were not observed, at least at this temperature and time scale. Additionally, the final structures after annealing, while not identical, did not show significant differences between the compositions. The final structures can be found in the supplementary materials (Fig. S5).
@@ IV Discussion
Our general-purpose interatomic potential developed for AlCrCuFeNi alloys enables computationally efficient and accurate computational investigations of various alloys. While the resulting tabGAP model does not explicitly include magnetic degrees of freedom, the use of magnetic DFT training data is key to enable simulations of multiphase alloys and phase transitions. This is evident from the Monte Carlo swapping MD simulations, where a dynamic transition from FCC to BCC was observed for the equiatomic composition. Magnetism has been studied experimentally recently in the AlCrCuFeNi alloy, where it was noted the alloys exhibit robust hard ferromagnetism [103]. Furthermore, L. Kush et al. have estimated the Curie temperature of the alloy between 341.7 to 404.9 K depending on the composition [104, 105]. It is also stated that below the Curie temperature the alloy has a very stable ferromagnetic order. This means that the assumption of the initial ferromagnetic order in the DFT training data is acceptable. However, at higher temperatures or due to compositional variations and ordering, the alloy may exhibit more complex magnetic behavior, limiting the transferability of the current model. In such cases a more accurate model that takes into account the explicit magnetic degree of freedom will be required. Even more strikingly, the MCMD simulations computationally support the VEC stability rule (see Table 1 for VECs), where for alloy compositions with VEC $\geq 8$ a single FCC phase is expected and for $6.87\leq$ VEC << 8 FCC and BCC coexistence is expected [80]. This rule has also found support in experiments [10, 106, 107]. However, the VEC rule might not be well suited for prediction of phases in non-equilibrium rapidly solidified HEAs [15], and BCC phases have been observed with AlCrCuFeNi alloys above the VEC criterion [95]. Furthermore, the addition of large amounts of Al causes significant lattice distortions (due to it having a large atom radius, see Table 1 for $\delta\%$), which makes the close-packed FCC less stable and thus the BCC lattice can accommodate the larger Al atoms more easily in it [108].